OpenStudy (anonymous):

Solve by taking the square roots of both sides. (x + 4)2 - 5 = 2

OpenStudy (anonymous):

Do you want steps on this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k first what do you think it is

OpenStudy (anonymous):

I don know its confusing to me

OpenStudy (anonymous):

Step 1: Simplify both sides of the equation. (x+4)(2)−5=2 Simplify: (Show steps) 2x+3=2 Step 2: Subtract 3 from both sides. 2x+3−3=2−3 2x=−1 Step 3: Divide both sides by 2. 2x = -1 --- --- 2 2 x= −1 2 x −1 ----- 2 ________________________________________ Answer: x= −1 2

OpenStudy (anonymous):

but it doesn't have that as a choice

OpenStudy (anonymous):

wHAT DO YOU HAVE

OpenStudy (anonymous):

caps

OpenStudy (anonymous):

what do you have for answers

OpenStudy (anonymous):

a.3,x=-11 b.4+7 squared ,x=-4+7 squared c. x=-4+7 squared , x=-4-7 squared d. x=-4+11 squared , x=-4-11 squared

OpenStudy (anonymous):

really need help with this though